Write an equation for an ellipse with vertices at (-2,-1) and (2,-7) and co-vertices at (-4,3) and (0,3)
Answer by lwsshak3(11628)

Write an equation for an ellipse with vertices at (-2,-1) and (2,-7) and co-vertices at (-4,3) and (0,3)
This is an ellipse with vertical major axis.
Its standard form of equation: (x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k)=(x,y) coordinates of center
y-coordinate of center=-4 (midpoint between -1 and -7)
x-coordinate of center=-2 (midpoint between 0 and -4
center:(-2,-4)
length of vertical major axis=6,(-1 to -7)=2a
a=3
a^2=9
..
length of minor axis=4,(-4 to 0)=2a
b=2
b^2=4
..
Equation:
(x+2)^2/4+(y+4)^2/9=1